OpenStudy (klepez):

Simplify (7t2 + 9) + (6t2 + 8).

OpenStudy (anonymous):

\[(7t ^{2}+9)+(6t ^{2}+8)\] \[=7t ^{2}+9+6t ^{2}+8\] \[=7t ^{2}+6t ^{2}+8+9\] \[=13t ^{2}+17\]

OpenStudy (anonymous):

1. I rewrote the expression 2. Can simply remove the parentheses because there are no negative signs to worry about 3. Put like terms together 4. Simplify by combining the like terms

OpenStudy (klepez):

thanks